# Latest activities and results from T2K

### v1

@@ 1 Introduction and physics goals of T2K
Figure 1: Unresolved questions about neutrinos and leptonic mixing: There are discrete options for the mass ordering and octant, and in only one of them (NO, $\theta_{23}<\frac{\pi}{2}$) would the leptons retain a generational (i.e. leading diagonal) structure. The matrix may also have a unitary form, leading to CP violation.
@@ 2 T2K-only analysis
T2K posterior $\theta_{23}<\frac{\pi}{2}$ $\theta_{23}>\frac{\pi}{2}$
@@
Figure 2: [L] Evolution of confidence intervals on the CP phase. The new interaction model and change to the reactor constraint both widen the intervals, slightly more than the additional data helps constrain them. [R] Posterior probabilities for the mass ordering ($\mathrm{sign}[\Delta m^{2}_{32}]$), and octant ($\mathrm{sign}[\theta_{32}-\frac{\pi}{2}]$) degeneracies, obtained using T2K data only.
@@ 3 T2K+SK analysis
[…]
@@
T2K+SK posterior $\theta_{23}<\frac{\pi}{2}$ $\theta_{23}>\frac{\pi}{2}$

### v3

@@ 1 Introduction and physics goals of T2K
Figure 1: Unresolved questions about neutrinos and leptonic mixing: There are discrete options for the mass ordering and octant, and in only one of them (NO, $\theta_{23}<\frac{\pi}{4}$) would the leptons retain a generational (i.e. leading diagonal) structure. The matrix may also have a unitary form, leading to CP violation.
@@ 2 T2K-only analysis
T2K posterior $\theta_{23}<\frac{\pi}{4}$ $\theta_{23}>\frac{\pi}{4}$
@@
Figure 2: [L] Evolution of confidence intervals on the CP phase. The new interaction model and change to the reactor constraint both widen the intervals, slightly more than the additional data helps constrain them. [R] Posterior probabilities for the mass ordering ($\mathrm{sign}[\Delta m^{2}_{32}]$), and octant ($\mathrm{sign}[\theta_{32}-\frac{\pi}{4}]$) degeneracies, obtained using T2K data only.
@@ 3 T2K+SK analysis
The combined T2K+SK analysis uses the complimentary strengths of the two experiments, and is a much deeper integration than the T2K+NOvA joint analysis, as shown in Figure 3. The two experiments share the SK detector, so the associated uncertainties are strongly correlated (although not entirely, because of the different flux). The interaction model is also shared, which is what allows T2K’s ND280 data to constrain SK’s cross-section uncertainties.
@@
T2K+SK posterior $\theta_{23}<\frac{\pi}{4}$ $\theta_{23}>\frac{\pi}{4}$